OpenStudy (anonymous):

Solve: cos2Ɵ = sin Ɵ

myininaya (myininaya):

\[\cos(2 \theta)-\sin(\theta)=0\] Hint put in terms of sine ------------------------- \[\cos(2 \theta)=\cos^2(\theta)- \sin^2(\theta)=(1-\sin^2(\theta))-\sin^2(\theta)=1-2\sin^2(\theta)\]

OpenStudy (anonymous):

is 1−2sin2(θ) the answer ?

myininaya (myininaya):

No

OpenStudy (anonymous):

how do i go from there?

myininaya (myininaya):

\[\cos(2 \theta)=1 - 2 \sin^2(\theta) \text{ is an identity}\] we need to use it in our problem

myininaya (myininaya):

\[1-2\sin^2(\theta)-\sin(\theta)=0\] you need to solve this for theta

myininaya (myininaya):

\[-2 \sin^2(\theta)-\sin(\theta)+1=0 \] Can you solve: \[-2u^2-u+1=0\] for u?

myininaya (myininaya):

if so the relationship btw the equations I put is \[u=\sin(\theta)\] so find what u is then replace u with sin(theta)

OpenStudy (anonymous):

the problem above −2u2−u+1=0 you just add u to the other side right?

myininaya (myininaya):

No solve that for u

myininaya (myininaya):

\[-2u^2-u+1=0\] Solve this for u

OpenStudy (anonymous):

u = √1/2

myininaya (myininaya):

You should had got two solutions.

OpenStudy (anonymous):

Square roots can be both negative and positive. Those would be the two solutions.

OpenStudy (anonymous):

oh i see

OpenStudy (anonymous):

so it's u= +-√1/2 ?

OpenStudy (anonymous):

Yes

myininaya (myininaya):

\[(-2u+1)(u+1)=0\]

OpenStudy (anonymous):

is it 2u + 1 = 0?

myininaya (myininaya):

You have a*b=0 => either a=0 , b=0 or both=0 so we have -2u+1=0 or u+1=0 Solve both of these linear equations.

OpenStudy (anonymous):

-2u+1 i knew it :|

OpenStudy (anonymous):

u = 1/2 and u = -1

myininaya (myininaya):

Write now remember our relationship \[u=\sin(\theta)\]

myininaya (myininaya):

So we have \[\sin(\theta)=\frac{1}{2} \text{ or } \sin(\theta)=-1\]

myininaya (myininaya):

Oops right* not write* sorry wrong right/write lol

OpenStudy (anonymous):

it's ok

OpenStudy (anonymous):

so I can plug in to the calculator : inverse sin (1/2) then

OpenStudy (anonymous):

after that multiply by pi/180 to get pi/6 right ?

OpenStudy (anonymous):

30 or Pi/6

myininaya (myininaya):

Well I would take off points if you used a calculator

OpenStudy (anonymous):

oh ok

myininaya (myininaya):

It doesn't give you all the solutions and plus sometimes it gives you an approximation

myininaya (myininaya):

depending on what calculator you use

myininaya (myininaya):

There is the unit circle though

OpenStudy (anonymous):

it's really confusing :|

OpenStudy (anonymous):

but if i use calculator then I'm fine with that :)

myininaya (myininaya):

when is sine, 1/2?

myininaya (myininaya):

brainshot gave you one value that satisfies the equation

OpenStudy (anonymous):

@myininaya Are you a teacher?

OpenStudy (anonymous):

Myininaya is the awsomest teacher of all time :D

OpenStudy (anonymous):

Ha yeah, but I meant in reality.

OpenStudy (anonymous):

Oh yes in reality.

OpenStudy (anonymous):

What do you teach?

OpenStudy (anonymous):

Calculus to undegrad students in a university.

OpenStudy (anonymous):

Lol, I am talking for her :P :D

myininaya (myininaya):

@brainshot3 I am. \[\sin(\theta)=\frac{1}{2}\] \[\theta=\frac{\pi}{6}+2 n \pi \text{ this } n 2 \pi \text{ part is the n times we go around the circle}\] But we also have another value in the interval [0,2pi] such that \[\sin(\theta)=\frac{1}{2}\] \[\theta=\frac{5 \pi}{6} +2 n \pi\] And by the way n is an integer

OpenStudy (anonymous):

and a phd student in mathematics and yet find time to help people here lol

OpenStudy (anonymous):

Wow! What do you plan to get a job in?

myininaya (myininaya):

You really need to familiarize yourself with the unit circle. It is your best friend in trig.

OpenStudy (anonymous):

thanks for the link!

myininaya (myininaya):

You still need to solve the other equation.

myininaya (myininaya):

\[\sin(\theta)=-1 \]

OpenStudy (anonymous):

i got cos^2 θ = 0 ?

OpenStudy (anonymous):

thanks for the link @FoolForMath

OpenStudy (anonymous):

\[ \sin(\theta)=-1=\sin (-\frac \pi 2) \implies \theta = n\pi +(-1)^n (-\frac \pi 2) \]

OpenStudy (anonymous):

Btw, Myin, I thought that definition of \(2n\pi \) is only for \( \cos \), am I missing something?

OpenStudy (anonymous):

This question is really hard for me, I will look over it again tonight thanks guys!

OpenStudy (anonymous):

Btw can you guys help me with another one ?

OpenStudy (anonymous):

I kinda in a rush :|

OpenStudy (anonymous):

\[ \sin(\theta)=-1=\sin (-\frac \pi 2) \implies \theta = n\pi +(-1)^n (-\frac \pi 2) \]\[= n\pi +(-1)^{n+1} \left( \frac \pi 2\right)\]